You are standing on a bathroom scale inside an elevator. Your weight (as measured on a bathroom scale under ordinary conditions) is 168.3 lb, but reading on the scale is 146.5 lb. What is the y-component of the acceleration?
Answer #1

consider the force on the scale

N - W = m a

where

  • N = 146.5 lb-f
  • W = 168.3 lb-f
  • m = W / g = ( 168.3 lb ) / ( 32,17 ft/s² ) = 5.232 lb

then

a = ( 146.5 lb-f - 168.3 lb-f ) / ( 5.232 lb )

a = -4.167 ft/s²